# ObjectAdd: Adding Objects into Image via a Training-Free Diffusion Modification Fashion

### v5

@@ Abstract
Index Terms:
Diffusion Model, Training-Free, Text to Image, Inpainting
@@ I Introduction
Recent years have seen significant advancements in transformer architecture [1], like CLIP [2], which have improved natural language embedding methods. Additionally, the development of diffusion models [3] has led to more accurate and diverse image generation. These advancements have jointly made text-to-image generation increasingly popular. Typically, a text-to-image model takes a description sentence as input and generates a description-aware image. Models such as DALL-E 3 [4] and Stable Diffusion [5] have been particularly noteworthy for their impressive performances. On the premise of these developments, many software platforms have been built to enable non-professionals to easily generate images from text, which has a significant impact on the art industry.
@@
One example is Midjourney, where users can create and share AI-generated images [6]. Another example is ChatGPT, of which the latest version allows users to request image generation [7]. Even classical software tools like Photoshop have improved their functionality by incorporating AI technology, enabling them to perform previously unimaginable tasks, such as changing the orientation of objects in photos [8]. These applications pave the way for the future of artificial general intelligence (AGI) as they can be operated solely through natural language commands. With the use of these tools, even individuals without any formal training can now create their own artwork, leading to a significant boost in the efficiency of the art industry. However, most text-to-image models rely solely on text prompts as inputs. This makes it challenging for users to convey spatial relationships between objects. Take the most famous large model, ChatGPT, for example. It requires users to describe all their requirements through chatting, which is not an easy task for non-professional personnel. Some approaches have been proposed to solve this problem, with ControlNet [9] possibly being one of the most famous models. ControlNet allows users to include canny edge or human pose as extra inputs, enabling control over the pose of humans or animals in the generated image. With the use of ControlNet, users can control the generated result in a more intuitive way, but it still requires users to decide the layout all at once before seeing any outcome result.
@@
In this paper, we propose ObjectAdd, a training-free method that allows users to add an object into a selected area of an image, while maintaining the rest content almost unchanged. Given an image prompt and object word, we start by analyzing the negative impact of self-attention mechanism in the text encoder, causing embedding confusion if conducting a simple text concatenation. Therefore, the first innovation of our ObjectAdd is to attempt concatenating the embedding outputs and amalgamate information from both texts while preventing them from interfering with each other. Then, we continue our second innovation of object-driven layout control in order to capture objects within the user-drawn area. ObjectAdd takes into consideration both latent- and attention-level information into the diffusion process of the edited image. For latent information injection, an object-oriented diffusion process is individually conducted, during which, the training-free backward guidance [10] is firstly adopted to strengthen the cross-attention score of added objects inside the user-drawn box, then the corresponding area in diffusion latent is transferred to that of edited image. As for attention information injection, we enhance the cross-attention map in the user-drawn area of edited image. By doing so in the early diffusion steps, we realize adding a high-quality object to the edited image.
@@ II-A Training-Free Control
Since large text-to-image models are trained with extensive datasets, they are generally demonstrated to embrace enough information for most tasks. Given that retraining a large model can be challenging and may lead to catastrophic forgetting [11], many researchers focus on training-free methods to control the model outputs. These methods fall into two main categories: forward and backward guidance. In our work, we combine both approaches.
@@
To modify the output of a pretrained model without retraining, a straightforward approach is to directly adjust the parameters without employing gradient descent. Since this approach doesn’t involve backward propagation, they are referred to as “forward guidance.” Most diffusion-based models incorporate attention maps, which contain semantic information [12]. Many forward guidance methods modify attention maps to improve their effectiveness [12, 13, 14]. The two main methods for editing attention maps are enhancing and swapping. Enhancement is often applied to the cross-attention map. This increases the value in the target area of the map, ensuring that objects are generated correctly [13]. Swapping, however, can be applied to both self- and cross-attention maps. In some studies [12], researchers swap parts of the cross-attention map to replace one object with another, while keeping the rest of the image unchanged. In other cases [14], parts of the self attention map from the original image are swapped with the editing output to maintain consistency.
@@
Although researchers sometimes expect to eliminate training, having a loss function is still an easy way to guide the model. Methods that require no training but utilize a loss function and gradient descent are known as “backward guidance.” These methods are commonly used to control the layout of generated images [10, 15, 16]. Compared to forward guidance, backward guidance imposes stricter constraints on object placement. It typically identifies object locations using cross-attention maps [10, 15, 16], and calculates loss based on user-marked boxes or areas. The loss value is then used to apply gradient descent to the latent, which is updated iteratively at each time step. This process continues until it reaches the maximum number of iterations or the loss falls below a certain threshold. Backward guidance is also effective for repositioning objects. For instance, in DragDiffusion [17], it is used to move selected objects in a specific direction, and user-drawn boxes are therefore no longer necessary.
@@ II-B Training-Dependent Image Editing
Though fine-tuning a large model could be difficult, an external model can be trained to assist the larger model. For instance, in ITI-Gen [18], researchers train an encoder to obtain distinctive token embeddings. This is done to accurately generate human faces that match the descriptive words in the prompt, such as skin tone. Another example is seen in GLIGEN [19], where an extra self-attention layer is trained and integrated into the pretrained model. This layer encodes and transmits location information to the model. The benefit of this method is its stability compared to training-free approaches. However, it requires a dataset tailored to its specific task, which can sometimes be challenging to obtain. InstructPix2Pix [20] could be the one that has a most similar goal to our approach. It allows users to specify their desired modifications to an input image through prompts. This model takes input image as conditional vector, as well as the instruction prompt indicated by user. The output of this approach is the edit version of input image, following the instruction. However, it still only allows prompts as input for specifying the requirements, and its results often fail to guarantee the consistency.
@@ II-C Decoupled Text Embedding
Most current popular text encoders include self-attention layers, implying that the encoding of a word can be influenced by other words within the input. This poses a challenge known as prompt concatenation. Simply combining two prompts and encoding them together often cause confusion. For instance, consider the two prompts “a cat” and “a dog” combined to form a new prompt “a cat a dog.” This combination lacks in-context meaning, making it challenging for the model to comprehend. The solution to this problem involves encoding the prompts separately and then concatenating their embeddings. The selection of separately encoded words may vary; in the case of ITI-GEN [18], category-related words are encoded individually, while in another paper [21], all nones are encoded individually. Both of the papers observe increase performance.
@@ II-D Diffusion Based Real Image Editing
Diffusion models have demonstrated significant capabilities in controllable image generation [12, 22]. There is a growing interest in using them for real image editing. Since diffusion models start with noise as input, the primary challenge to edit real images is to convert the input real image into noise in such a way that the noise can be reconstructed into the original image after the denoising steps. This conversion process is called inversion. The most commonly used method today is DDIM inversion [23], which inverts each denoising step by adding the predicted noise to the latent representation. Several approaches have been proposed to improve the quality of inversion. In [24], the researchers suggested that inversion could be considered a fixed-point iteration problem and could be solved using iterative updating methods. ReNoise [25] addresses the inversion problem from a similar perspective, but introduces minor biases in each step to enhance the edit-ability of the inversion results.
Once the real image has been inverted to noise, it can be used as the input for the diffusion model. Various methods can then be applied to control the generation process, thereby editing the real image. MasaCtrl [22] is one example. In that paper, researchers replace information in the editing branch with the corresponding information in the reconstruction branch to maintain consistency while using new prompts as conditions for generation. This approach allows editing original image, applying modifications while retaining most of the original information.
@@ III-A Preliminaries
Herein, $H$, $W$ and $C$ respectively indicate image height, width and channel number. Inside the stable diffusion model $F$, there are some intermediate representations that are the focus in our method: (1) Text embedding $\mathcal{E}_{\mathcal{P}}\in\mathbb{R}^{N\times D}$, usually derived from a CLIP model [2], where $N$ represents the maximum token sequence length in the text encoder and $D$ denotes the embedding dimension; (2) The $t$-th step latent denoised representation $\mathcal{I}^{t}\in\mathbb{R}^{H\times W\times C}$; (3) The $\gamma$-layer cross-attention map at the $t$-th step $\mathcal{C}^{\gamma t}\in\mathbb{R}^{N\times(H_{\gamma}\cdot W_{\gamma})}$, with $\mathcal{C}^{\gamma t}_{i,:}\in\mathbb{R}^{H_{\gamma}\cdot W_{\gamma}}$ indicating the attentive score of the $\gamma$-th layer output to the $i$-th token. For ease of the following representation, we reshape $C^{{\gamma}t}_{i,:}$ into $H_{\gamma}\times W_{\gamma}$. Notice these symbols will be revised accordingly in the following sections, in order to adapt to outputs of different prompts. For example, $\mathcal{C}^{{\gamma}t}$ will be revised as $\bar{\mathcal{C}}^{{\gamma}t}$ for adapting to Eq. (2) and $\tilde{\mathcal{C}}^{{\gamma}t}$ for adapting to Eq. (4).
Our intend to edit on image $\mathcal{I}^{0}$ to derive a user-expected version ¯$\bar{\mathcal{I}}^{0}$ under two application-oriented settings: 1) a user-drawn box in image $\mathcal{I}^{0}$ and 2) a user-specified word-level prompt $\mathcal{W}$ describing an object. The box can be characterised as a binary mask $\mathcal{M}\in\{0,1\}^{H\times W}$ where $\mathcal{M}_{i,j}=1$ if pixel $\mathcal{I}^{0}_{i,j}$ falls into the drawn box, and $\mathcal{M}_{i,j}=0$ else. Note that, the $\mathcal{M}$ is downsampled into $\mathcal{M}^{\gamma}\in\mathbb{R}^{H_{\gamma}\times W_{\gamma}}$, for adapting to the $\gamma$-th layer image shape.
The current edited image ¯$\bar{\mathcal{I}}^{0}$ is featured with: 1) the word $\mathcal{W}$ described object inside the box area; 2) exact content with $\mathcal{I}^{0}$ outside the box area; 3) flawless fusion between the two areas. The evolution process of ¯$\bar{\mathcal{I}}^{0}$ can be described as:
¯$\bar{\mathcal{I}}^{0}=F(\mathcal{I}^{T};\{\mathcal{P},\mathcal{W}\}),\qquad\mathcal{I}^{T}\sim\mathcal{N}(0,\mathbf{I}).$ (2)
@@ III-B Text Embedding Coalesce
We first deal with the coalesce of prompts $\mathcal{P}$ and $\mathcal{W}$ in Eq. (2). As specified in Sec. II-C, the self-attention mechanism in text encoder causes embedding confusion if simply concatenating two prompts together $concat(\mathcal{P},\mathcal{W})$. Taking $\mathcal{P}$ = “A woman wearing glasses” and $\mathcal{W}$ = “A hat” as a simple example, the concatenation results in prompt “A woman wearing glasses A hat” tokenized as (“[CLS],” “a,” “woman,” “wearing,” “glasses,” “a,” “hat,” “[SEP]”) where “[CLS]” and “[SEP]” respectively denote the start and end tokens. For self-attention without causal masks, these tokens are mutually attentioned therefore causing different embeddings of “a,” “woman,” “wearing,” “glasses,” “a,” and “hat,” compared to individually embed prompts $\mathcal{P}$ (“[CLS],” “a,” “woman,” “wearing,” “glasses,” “[SEP]”) and $\mathcal{W}$ (“[CLS],” “a,” “hat,” “[SEP]”). Similarly, masked self-attention causes different embeddings of “a” and “hat.” Most current diffusion models deploy a masking-free CLIP-style text encoder [2], which contradicts the intention in this paper that users add an object into a selected area of an image, while maintaining the rest content unchanged.
@@ III-C Object-Driven Layout Control
We continue addressing the second issue of object-driven layout control in Eq. (2). To generate the word $\mathcal{W}$ described object within the user-drawn box area, our ObjectAdd injects both latent- and attention-level information into the diffusion process of the edited image ¯$\bar{\mathcal{I}}^{0}$. Below, we denote $k$ as the token sequence id of added object after the text embedding coalesce.
@@ III-C1 Latent Information Injection
~$\tilde{\mathcal{I}}^{0}=F(\mathcal{I}_{T};\mathcal{W}),\qquad\mathcal{I}^{T}\sim\mathcal{N}(0,\mathbf{I}).$ (4)
To control $\mathcal{W}$ described object accessing the user-drawn box area, we choose the training-free backward guidance method [10], basic principle of which is to restrict the size and location of the object. Recall the cross-attention map $\tilde{C}^{{\gamma}t}_{k,:}\in\mathbb{R}^{H_{\gamma}\times W_{\gamma}}$ represents attentiveness of the added object to the $\gamma$-th layer image. The learning objective can be generally formulated as:
@@
Further, the backward propagation process is conducted to guide the latent denoised representation ~$\tilde{\mathcal{I}}^{t}$ in Eq. (4) as:
~$\tilde{\mathcal{I}}^{t}\leftarrow\tilde{\mathcal{I}}^{t}-\eta\nabla_{\tilde{\mathcal{I}}^{t}}\sum_{\gamma\in\Gamma}E\left(\tilde{\mathcal{C}}^{{\gamma}t},\mathcal{M}^{\gamma},k\right),$ (6)
which iterates several times. The $\Gamma$ represents all chosen layers of the cross-attention maps and $\eta$ is a learning rate. By shifting the cross-attention focus, it has been well demonstrated [10] that the word $\mathcal{W}$ described object can be mostly confined to the user-drawn box area of latent ~$\tilde{\mathcal{I}}^{t}$.
The goal of this paper is to add an object into image ¯$\bar{\mathcal{I}}^{0}$, prompted by text embedding coalesce $\mathcal{E}_{\mathcal{P},\mathcal{W}}$. Thus, we further inject the above updated latent ~$\tilde{\mathcal{I}}^{t}$ into the intended latent ¯$\bar{\mathcal{I}}^{t}$ at the $t$-th time step as:
¯$\bar{\mathcal{I}}^{t}=(1-\mathcal{M})\odot\bar{\mathcal{I}}^{t}+\mathcal{M}\odot\tilde{\mathcal{I}}^{t},$ (7)
@@
Figure 2: Generation comparison under settings of $\mathcal{P}$ = “A woman wearing glasses” and $\mathcal{W}$ = “A hat”: (a) $\mathcal{P}$ prompted image $\mathcal{I}^{0}$; (b) the coalesce $\mathcal{E}_{\{\mathcal{P},\mathcal{W}\}}$ prompted image ¯$\bar{\mathcal{I}}^{0}$ after object-driven layout control; (c) ¯$\bar{\mathcal{I}}^{0}$ with prompted image inpainting.

### v6

@@ I Introduction
Recent years have seen significant advancements in transformer architecture [vaswani2017attention], like CLIP [radford2021learning], which have improved natural language embedding methods. Additionally, the development of diffusion models [ho2020denoising] has led to more accurate and diverse image generation. These advancements have jointly made text-to-image generation increasingly popular. Typically, a text-to-image model takes a description sentence as input and generates a description-aware image. Models such as DALL-E 3 [betker2023improving] and Stable Diffusion [rombach2022high] have been particularly noteworthy for their impressive performances. On the premise of these developments, many software platforms have been built to enable non-professionals to easily generate images from text, which has a significant impact on the art industry.
@@
One example is Midjourney, where users can create and share AI-generated images [midjourney]. Another example is ChatGPT, of which the latest version allows users to request image generation [chatGPT]. Even classical software tools like Photoshop have improved their functionality by incorporating AI technology, enabling them to perform previously unimaginable tasks, such as changing the orientation of objects in photos [photoshop]. These applications pave the way for the future of artificial general intelligence (AGI) as they can be operated solely through natural language commands. With the use of these tools, even individuals without any formal training can now create their own artwork, leading to a significant boost in the efficiency of the art industry. However, most text-to-image models rely solely on text prompts as inputs. This makes it challenging for users to convey spatial relationships between objects. Take the most famous large model, ChatGPT, for example. It requires users to describe all their requirements through chatting, which is not an easy task for non-professional personnel. Some approaches have been proposed to solve this problem, with ControlNet [zhang2023adding] possibly being one of the most famous models. ControlNet allows users to include canny edge or human pose as extra inputs, enabling control over the pose of humans or animals in the generated image. With the use of ControlNet, users can control the generated result in a more intuitive way, but it still requires users to decide the layout all at once before seeing any outcome result.
@@
In this paper, we propose ObjectAdd, a training-free method that allows users to add an object into a selected area of an image, while maintaining the rest content almost unchanged. Given an image prompt and object word, we start by analyzing the negative impact of self-attention mechanism in the text encoder, causing embedding confusion if conducting a simple text concatenation. Therefore, the first innovation of our ObjectAdd is to attempt concatenating the embedding outputs and amalgamate information from both texts while preventing them from interfering with each other. Then, we continue our second innovation of object-driven layout control in order to capture objects within the user-drawn area. ObjectAdd takes into consideration both latent- and attention-level information into the diffusion process of the edited image. For latent information injection, an object-oriented diffusion process is individually conducted, during which, the training-free backward guidance [chen2023trainingfree] is firstly adopted to strengthen the cross-attention score of added objects inside the user-drawn box, then the corresponding area in diffusion latent is transferred to that of edited image. As for attention information injection, we enhance the cross-attention map in the user-drawn area of edited image. By doing so in the early diffusion steps, we realize adding a high-quality object to the edited image.
@@ II-A Training-Free Control
Since large text-to-image models are trained with extensive datasets, they are generally demonstrated to embrace enough information for most tasks. Given that retraining a large model can be challenging and may lead to catastrophic forgetting [zhai2024investigating], many researchers focus on training-free methods to control the model outputs. These methods fall into two main categories: forward and backward guidance. In our work, we combine both approaches.
@@
To modify the output of a pretrained model without retraining, a straightforward approach is to directly adjust the parameters without employing gradient descent. Since this approach doesn’t involve backward propagation, they are referred to as “forward guidance.” Most diffusion-based models incorporate attention maps, which contain semantic information [hertz2022prompt]. Many forward guidance methods modify attention maps to improve their effectiveness [hertz2022prompt, densediffusion, patashnik2023localizing]. The two main methods for editing attention maps are enhancing and swapping. Enhancement is often applied to the cross-attention map. This increases the value in the target area of the map, ensuring that objects are generated correctly [densediffusion]. Swapping, however, can be applied to both self- and cross-attention maps. In some studies [hertz2022prompt], researchers swap parts of the cross-attention map to replace one object with another, while keeping the rest of the image unchanged. In other cases [patashnik2023localizing], parts of the self attention map from the original image are swapped with the editing output to maintain consistency.
@@
Although researchers sometimes expect to eliminate training, having a loss function is still an easy way to guide the model. Methods that require no training but utilize a loss function and gradient descent are known as “backward guidance.” These methods are commonly used to control the layout of generated images [chen2023trainingfree, chefer2023attend, xie2023boxdiff]. Compared to forward guidance, backward guidance imposes stricter constraints on object placement. It typically identifies object locations using cross-attention maps [chen2023trainingfree, chefer2023attend, xie2023boxdiff], and calculates loss based on user-marked boxes or areas. The loss value is then used to apply gradient descent to the latent, which is updated iteratively at each time step. This process continues until it reaches the maximum number of iterations or the loss falls below a certain threshold. Backward guidance is also effective for repositioning objects. For instance, in DragDiffusion [shi2023dragdiffusion], it is used to move selected objects in a specific direction, and user-drawn boxes are therefore no longer necessary.
@@ II-B Training-Dependent Image Editing
Though fine-tuning a large model could be difficult, an external model can be trained to assist the larger model. For instance, in ITI-Gen [zhang2023iti], researchers train an encoder to obtain distinctive token embeddings. This is done to accurately generate human faces that match the descriptive words in the prompt, such as skin tone. Another example is seen in GLIGEN [li2023gligen], where an extra self-attention layer is trained and integrated into the pretrained model. This layer encodes and transmits location information to the model. The benefit of this method is its stability compared to training-free approaches. However, it requires a dataset tailored to its specific task, which can sometimes be challenging to obtain. InstructPix2Pix [brooks2023instructpix2pix] could be the one that has a most similar goal to our approach. It allows users to specify their desired modifications to an input image through prompts. This model takes input image as conditional vector, as well as the instruction prompt indicated by user. The output of this approach is the edit version of input image, following the instruction. However, it still only allows prompts as input for specifying the requirements, and its results often fail to guarantee the consistency.
@@ II-C Decoupled Text Embedding
Most current popular text encoders include self-attention layers, implying that the encoding of a word can be influenced by other words within the input. This poses a challenge known as prompt concatenation. Simply combining two prompts and encoding them together often cause confusion. For instance, consider the two prompts “a cat” and “a dog” combined to form a new prompt “a cat a dog.” This combination lacks in-context meaning, making it challenging for the model to comprehend. The solution to this problem involves encoding the prompts separately and then concatenating their embeddings. The selection of separately encoded words may vary; in the case of ITI-GEN [zhang2023iti], category-related words are encoded individually, while in another paper [feng2022training], all nones are encoded individually. Both of the papers observe increase performance.
@@ II-D Diffusion Based Real Image Editing
Diffusion models have demonstrated significant capabilities in controllable image generation [hertz2022prompt, cao_2023_masactrl]. There is a growing interest in using them for real image editing. Since diffusion models start with noise as input, the primary challenge to edit real images is to convert the input real image into noise in such a way that the noise can be reconstructed into the original image after the denoising steps. This conversion process is called inversion. The most commonly used method today is DDIM inversion [couairon2023diffedit], which inverts each denoising step by adding the predicted noise to the latent representation. Several approaches have been proposed to improve the quality of inversion. In [pan2023effective], the researchers suggested that inversion could be considered a fixed-point iteration problem and could be solved using iterative updating methods. ReNoise [garibi2024renoise] addresses the inversion problem from a similar perspective, but introduces minor biases in each step to enhance the edit-ability of the inversion results.
Once the real image has been inverted to noise, it can be used as the input for the diffusion model. Various methods can then be applied to control the generation process, thereby editing the real image. MasaCtrl [cao_2023_masactrl] is one example. In that paper, researchers replace information in the editing branch with the corresponding information in the reconstruction branch to maintain consistency while using new prompts as conditions for generation. This approach allows editing original image, applying modifications while retaining most of the original information.
@@ III-A Preliminaries
Herein, $H$, $W$ and $C$ respectively indicate image height, width and channel number. Inside the stable diffusion model $F$, there are some intermediate representations that are the focus in our method: (1) Text embedding $\mathcal{E}_{\mathcal{P}}\in\mathbb{R}^{N\times D}$, usually derived from a CLIP model [radford2021learning], where $N$ represents the maximum token sequence length in the text encoder and $D$ denotes the embedding dimension; (2) The $t$-th step latent denoised representation $\mathcal{I}^{t}\in\mathbb{R}^{H\times W\times C}$; (3) The $\gamma$-layer cross-attention map at the $t$-th step $\mathcal{C}^{\gamma t}\in\mathbb{R}^{N\times(H_{\gamma}\cdot W_{\gamma})}$, with $\mathcal{C}^{\gamma t}_{i,:}\in\mathbb{R}^{H_{\gamma}\cdot W_{\gamma}}$ indicating the attentive score of the $\gamma$-th layer output to the $i$-th token. For ease of the following representation, we reshape $C^{{\gamma}t}_{i,:}$ into $H_{\gamma}\times W_{\gamma}$. Notice these symbols will be revised accordingly in the following sections, in order to adapt to outputs of different prompts. For example, $\mathcal{C}^{{\gamma}t}$ will be revised as $\bar{\mathcal{C}}^{{\gamma}t}$ for adapting to Eq. (2) and $\tilde{\mathcal{C}}^{{\gamma}t}$ for adapting to Eq. (4).
Our intend to edit on image $\mathcal{I}^{0}$ to derive a user-expected version $\bar{\mathcal{I}}^{0}$ under two application-oriented settings: 1) a user-drawn box in image $\mathcal{I}^{0}$ and 2) a user-specified word-level prompt $\mathcal{W}$ describing an object. The box can be characterised as a binary mask $\mathcal{M}\in\{0,1\}^{H\times W}$ where $\mathcal{M}_{i,j}=1$ if pixel $\mathcal{I}^{0}_{i,j}$ falls into the drawn box, and $\mathcal{M}_{i,j}=0$ else. Note that, the $\mathcal{M}$ is downsampled into $\mathcal{M}^{\gamma}\in\mathbb{R}^{H_{\gamma}\times W_{\gamma}}$, for adapting to the $\gamma$-th layer image shape.
The current edited image $\bar{\mathcal{I}}^{0}$ is featured with: 1) the word $\mathcal{W}$ described object inside the box area; 2) exact content with $\mathcal{I}^{0}$ outside the box area; 3) flawless fusion between the two areas. The evolution process of $\bar{\mathcal{I}}^{0}$ can be described as:
$\bar{\mathcal{I}}^{0}=F(\mathcal{I}^{T};\{\mathcal{P},\mathcal{W}\}),\qquad\mathcal{I}^{T}\sim\mathcal{N}(0,\mathbf{I}).$ (2)
@@ III-B Text Embedding Coalesce
We first deal with the coalesce of prompts $\mathcal{P}$ and $\mathcal{W}$ in Eq. (2). As specified in Sec. II-C, the self-attention mechanism in text encoder causes embedding confusion if simply concatenating two prompts together $concat(\mathcal{P},\mathcal{W})$. Taking $\mathcal{P}$ = “A woman wearing glasses” and $\mathcal{W}$ = “A hat” as a simple example, the concatenation results in prompt “A woman wearing glasses A hat” tokenized as (“[CLS],” “a,” “woman,” “wearing,” “glasses,” “a,” “hat,” “[SEP]”) where “[CLS]” and “[SEP]” respectively denote the start and end tokens. For self-attention without causal masks, these tokens are mutually attentioned therefore causing different embeddings of “a,” “woman,” “wearing,” “glasses,” “a,” and “hat,” compared to individually embed prompts $\mathcal{P}$ (“[CLS],” “a,” “woman,” “wearing,” “glasses,” “[SEP]”) and $\mathcal{W}$ (“[CLS],” “a,” “hat,” “[SEP]”). Similarly, masked self-attention causes different embeddings of “a” and “hat.” Most current diffusion models deploy a masking-free CLIP-style text encoder [radford2021learning], which contradicts the intention in this paper that users add an object into a selected area of an image, while maintaining the rest content unchanged.
@@ III-C Object-Driven Layout Control
We continue addressing the second issue of object-driven layout control in Eq. (2). To generate the word $\mathcal{W}$ described object within the user-drawn box area, our ObjectAdd injects both latent- and attention-level information into the diffusion process of the edited image $\bar{\mathcal{I}}^{0}$. Below, we denote $k$ as the token sequence id of added object after the text embedding coalesce.
@@ III-C1 Latent Information Injection
$\tilde{\mathcal{I}}^{0}=F(\mathcal{I}_{T};\mathcal{W}),\qquad\mathcal{I}^{T}\sim\mathcal{N}(0,\mathbf{I}).$ (4)
To control $\mathcal{W}$ described object accessing the user-drawn box area, we choose the training-free backward guidance method [chen2023trainingfree], basic principle of which is to restrict the size and location of the object. Recall the cross-attention map $\tilde{C}^{{\gamma}t}_{k,:}\in\mathbb{R}^{H_{\gamma}\times W_{\gamma}}$ represents attentiveness of the added object to the $\gamma$-th layer image. The learning objective can be generally formulated as:
@@
Further, the backward propagation process is conducted to guide the latent denoised representation $\tilde{\mathcal{I}}^{t}$ in Eq. (4) as:
$\tilde{\mathcal{I}}^{t}\leftarrow\tilde{\mathcal{I}}^{t}-\eta\nabla_{\tilde{\mathcal{I}}^{t}}\sum_{\gamma\in\Gamma}E\left(\tilde{\mathcal{C}}^{{\gamma}t},\mathcal{M}^{\gamma},k\right),$ (6)
which iterates several times. The $\Gamma$ represents all chosen layers of the cross-attention maps and $\eta$ is a learning rate. By shifting the cross-attention focus, it has been well demonstrated [chen2023trainingfree] that the word $\mathcal{W}$ described object can be mostly confined to the user-drawn box area of latent $\tilde{\mathcal{I}}^{t}$.
The goal of this paper is to add an object into image $\bar{\mathcal{I}}^{0}$, prompted by text embedding coalesce $\mathcal{E}_{\mathcal{P},\mathcal{W}}$. Thus, we further inject the above updated latent $\tilde{\mathcal{I}}^{t}$ into the intended latent $\bar{\mathcal{I}}^{t}$ at the $t$-th time step as:
$\bar{\mathcal{I}}^{t}=(1-\mathcal{M})\odot\bar{\mathcal{I}}^{t}+\mathcal{M}\odot\tilde{\mathcal{I}}^{t},$ (7)
@@
Figure 2: Generation comparison under settings of $\mathcal{P}$ = “A woman wearing glasses” and $\mathcal{W}$ = “A hat”: (a) $\mathcal{P}$ prompted image $\mathcal{I}^{0}$; (b) the coalesce $\mathcal{E}_{\{\mathcal{P},\mathcal{W}\}}$ prompted image $\bar{\mathcal{I}}^{0}$ after object-driven layout control; (c) $\bar{\mathcal{I}}^{0}$ with prompted image inpainting.